OpenStudy (anonymous):

for the geometric sequence 1 2 4 find a8

OpenStudy (anonymous):

Common ratio,\[r=2\] First term,\[a_1=1\] Good so far?

OpenStudy (anonymous):

ogthot, please reply. Say something. Ask if you don't get something....

OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

SO you don't understand what I posted. Okay, so you don't understand anything, or just a part of my post?

OpenStudy (anonymous):

i just need the answer

OpenStudy (anonymous):

According to what I read from the Code of Conduct, this site prohibits straight forward answers. I can tutor you, and help you understand better...

OpenStudy (anonymous):

Getting straight answers won't get you anywhere. It will be quick, trust me. Just follow with me.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

Sure...

OpenStudy (anonymous):

So, do you understand what I mean when I say that r=2 (r is the common ratio) ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Yes, and we are also given that the first term, is 1.

OpenStudy (anonymous):

There is a formula, for any geometric sequence. \[a_n=a_1\times r^{n-1}\]

OpenStudy (anonymous):

DO you see the formula?

OpenStudy (anonymous):

i think would help a8=1(2)^8-1

OpenStudy (anonymous):

no, it wouldn't. You are just spoon feeding, and I want the user to grasp the process and the understanding of how to do these problems. It would help to get the answer, but not the learning.... don't mean to be rude.

OpenStudy (anonymous):

Okay, ogthot, again, there is a formula for geometric sequence.. \[a_n=a_1\times r^{n-1}\] where a.n is the Nth term, and r to the power (n-1) is the number of "jumps" that you would need to get to the nth term.

OpenStudy (anonymous):

Starting from the first term, you would have to multiply it times 3 to get the 4th term, and times 4 to get the 5th term, and times 5 to get the n6th term, and on... okay?

OpenStudy (anonymous):

(by multiply it, I mean multiply the first term times the common ratio)

OpenStudy (anonymous):